OpenStudy (anonymous):

(12x^3-8x^2+25x-7)divided by(6x+1) need help please :)

OpenStudy (anonymous):

you can use either the long division or synthetic division.

OpenStudy (anonymous):

http://www.purplemath.com/modules/polydiv2.htm this is my answer, but in dividing polynomials, it is needed to check it many times. \[2x ^{2}-\frac{ 5 }{ 3 }x+\frac{ 40 }{ 9 } remainder -11\frac{ 4 }{ 9 }\]